# Generalisation Bounds of Zero-Shot Economic Forecasting using Time Series Foundation Models

### v1

@@ Abstract
This study investigates zero-shot forecasting capabilities of Time Series Foundation Models (TSFMs) for macroeconomic indicators. We apply TSFMs to forecasting economic indicators under univariate conditions, bypassing the need for train bespoke econometric models using and extensive training datasets. Our experiments were conducted on a case study dataset, without additional customisation. We rigorously back-tested three state-of-the-art TSFMs (Chronos, TimeGPT and Moirai) under data-scarce conditions and structural breaks. Our results demonstrate that appropriately engineered TSFMs can internalise rich economic dynamics, accommodate regime shifts, and deliver well-behaved uncertainty estimates out of the box, while matching state-of-the-art multivariate models on this domain. Our findings suggest that, without any fine-tuning, TSFMs can match or exceed classical models during stable economic conditions. However, they are vulnerable to degradation in performances during periods of rapid shocks. The findings offer guidance to practitioners on when zero-shot deployments are viable for macroeconomic monitoring and strategic planning.
Keywords GDP forecasting, time series foundation models, time series forecasting, zero-shot forecasting
@@ 1 Introduction
Timely and reliable forecasts of gross domestic product (GDP) are essential inputs for policy, finance, and corporate planning worldwide. Governments embed medium-term growth scenarios in budget frameworks, debt-sustainability analyses, and multi-year spending envelopes [1]. Prudential regulators feed GDP paths into system-wide stress tests to gauge the resilience of banks and insurers under adverse macroeconomic conditions [2]. Investors rebalance portfolios when growth expectations shift, treating GDP as a succinct proxy for business-cycle momentum [3]. Multinational firms align production schedules, inventories, and staffing levels with headline and sectoral GDP projections [4], while credit-rating agencies and development institutions incorporate forward-looking GDP assumptions into sovereign-risk assessments and concessional-finance formulas [5].
@@
However, producing dependable forecasts remains challenging: structural breaks, measurement errors, and sudden shocks can quickly erode model performance, thereby motivating continuous innovation, from classical econometric combinations to modern machine-learning approaches [6]. Forecasting GDP accurately across countries presents significant challenges. Firstly, early national accounts releases are frequently revised, sometimes to the extent of reversing the sign of reported growth. This means models must contend with evolving truths rather than a fixed target [7]. Secondly, structural changes, such as those triggered by commodity-price super-cycles, financial crises, natural disasters, and global pandemics, can abruptly disrupt historical relationships and invalidate previously stable parameters [8, 9, 10, 11]. Lastly, sector-level GDP series display diverse seasonal patterns and react unevenly to external-demand shocks, hindering the transfer of information from one industry to another [12].
@@
For decades, classical time-series models, such as autoregressive integrated moving-average (ARIMA) frameworks [13] and vector autoregressions (VARs) [14], have been used as forecasting tools for this domain due to their transparency, tractability, and ease of re-estimation. However, their core assumptions of linearity and stable parameters rarely hold in practice, and sudden shifts in policy regimes and technology shocks can all invalidate coefficients calibrated on historical data, causing forecast accuracy to deteriorate rapidly as the horizon lengthens [15]. Empirical surveys have shown that beyond a few quarters ahead, even well-specified and rich multivariate systems seldom outperform simple persistence or random-walk benchmarks [16]. This challenge is further compounded by the constant moving of the goalposts in the form of data revisions, where preliminary GDP releases and other key indicators are often substantially updated, meaning models trained on early vintages chase a moving target and deliver the least reliability precisely when decision-makers most need clarity [17]. Recent studies have explored more sophisticated modelling approaches such as mixed-frequency factor and Bayesian models which integrate hundreds of monthly indicators to sharpen nowcasts [18], as well as tree-based ensembles and hybrid neural networks capable of discovering non-linear interactions [19]. However, the more recent modelling approaches are accompanied by high overheads in terms of expertise and computation for modest returns in improved accuracies, while being dependent on provision and access to real-time multivariate inputs [20].
@@
Recent advancements in AI have introduced a new class of forecasting tools, namely Time Series Foundation Models (TSFMs), with the potential of mitigating some of the challenges inherent in macroeconomic forecasting [21]. Having been trained on vast arrays of various time series, these models have the capacity to transfer the learned patterns from one domain to another without domain-specific re-training or fine-tuning, thus significantly reducing the modelling resource overheads [22]. These large pre-trained TSFMs like Amazon’s Chronos [23], Salesforce’s Moirai [24], and Nixtla’s TimeGPT [25] treat numeric sequences as language tokens and leverage transformer backbones trained extensive datasets. TimeGPT offers a “plug-and-play” API capabilities providing zero-shot forecasts without local fine-tuning, while Moirai extends this paradigm to multivariate forecasting with exogenous covariates. Models like Chronos have been used in literature on domains such as electricity, traffic, and retail data, where they have reported better performances than tuned statistical baselines with minimal feature engineering. Given the largely unexplored capabilities of these models in macroeconomic context, our study seeks to explore to what degree the current cutting-edge TSFMs can generalise across data vintages, sectors, and structural breaks in economic datasets, under the “out-of-the-box”, zero-shot settings. For our study, we used macroeconomic data from New Zealand as a case study, which presents an unusually demanding testbed, given the vulnerability of its small economy to external factors, as well as its exposure to commodity-price cycles, natural disasters, and external-demand shocks [26], together with the frequent revisions and thus the uncertainty of its national accounts estimates [27].
@@
While acknowledging that optimal macroeconomic forecasting ideally requires the integration of rich, real-time features capturing key economic indicators as model inputs, this study adopts a distinct approach. We focus specifically on evaluating the pure zero-shot capabilities of TSFMs without incorporating such external features or domain-specific covariates. This deliberate choice is made to establish fundamental performance limits and explore the inherent forecasting potential of these models based solely on their pre-training, thereby establishing baselines of their transfer abilities, given the current paucity of research applying TSFMs, particularly in a zero-shot manner to the economic domain. The contributions and novelty of this work can be summarised as:
  • Empirical benchmark: we provide the zero-shot evaluation of leading TSFMs (Chronos, Moirai, TimeGPT) against classical econometric baselines from the Reserve Bank of New Zealand (RBNZ) forecasts. This evaluation covers New Zealand’s national GDP and sectoral industries.
  • Performance measurement: we demonstrate that TSFMs outperform other classical methods across various horizons and thus we establish their utility under certain conditions.
  • Operational guidance: we offer actionable insights for policy analysts by mapping the boundary conditions under which zero-shot TSFMs serve as low-maintenance forecasting tools for humans or economists. We also identify scenarios where lightweight classical models remain preferable.
2 Related work
@@ 2.1 Forecasting Difficulty for Macroeconomic Indicators
Forecast accuracy for macroeconomic aggregates is fundamentally constrained by low signal-to-noise ratios. A long tradition of forecast-evaluation studies shows that, once the horizon stretches beyond the now-cast and the subsequent quarter, point predictions of real GDP growth seldom beat a naive random walk, let alone a purely random-direction guess, when measured against out-of-sample performance [16]. Persistence forecasts, which simply carry forward the latest observed growth rate, provide a standard benchmark over which, even multivariate econometric systems rarely improve upon this baseline after the first step in the forecast horizon [28]. Even median private-sector projections at a four-quarter horizon exhibit RMSEs statistically indistinguishable from persistence [29].
@@
To complement algorithmic advances, researchers now emphasise the timing of data arrival. Mixed-frequency and real-time approaches integrate high-frequency indicators, electronic-card transactions, and daily mobility into quarterly GDP nowcasts [35], providing policymakers near-instant feedback during shocks [36, 37]. Model builders also adapt specifications to the country context. In open economies like New Zealand, global commodity prices, foreign demand, and idiosyncratic domestic cycles jointly shape growth dynamics; assessing how well models internalize these influences remains an active line of inquiry [38].
@@ 2.2 Modern and Emerging Forecasting Approaches
Building on these statistical platforms, institution-specific suites provide operational nowcasts. The Federal Reserve Bank of New York’s medium-scale DSGE integrates theory-consistent shocks with factor information for policy analysis [45], whereas the Atlanta Fed’s GDPNow decomposes each GDP sub-aggregate via bridge equations and updates almost daily, offering a transparent, additive view of U.S. growth [46]. More recently, attention has shifted to machine learning and hybrid methods that relax linearity and exploit high-dimensional features. Ensemble trees (random forests, gradient boosting) already outperform factor and penalised-regression baselines on Dutch GDP nowcasts [44]. Deep neural networks, especially Long Short-Term Memory (LSTM) models, capture nonlinear temporal dependencies and have beaten ARIMA benchmarks in volatile settings [47]. Hybrid ensembles push further by blending economic structure with ML flexibility: weighting forecasts from a time-varying-coefficient DFM and a recurrent neural network by inverse MSE reduces U.S. GDP errors beyond either model alone, while broader model-averaging strategies remain popular for error reduction [48]. A clear progression from linear autoregressions through factor-based systems and institutional nowcasting dashboards to data-hungry ML hybrids each stage addressing limitations exposed by the last and setting the stage for evaluating emerging foundation-model approaches.
@@ 2.3 Zero-Shot TSFMs in Economic Forecasting
Early evidence that transformers can act as generic sequence learners came from the Frozen Pre-trained Transformer (FPT) experiment, where a language model backbone was kept entirely frozen but fine-tuned competitively across diverse time-series tasks, showing that self-attention can operate as a domain-agnostic computation engine [49]. Follow-up work systematises this line of research by mapping architectures, pre-training objectives, adaptation strategies, and data modalities [50]. While architectural variants such as encoder–decoder hybrids [51], sparse-attention blocks [52], and decomposition-style residual paths [53] now dominate the design space, none have yet been stress-tested on long-horizon macroeconomic indicators that mix low-frequency trends with regime shifts.
@@
TSFMs are based on the transformer architecture, shown in Figure 1. TSFMs aim for broad transferability with little or no task-specific fine-tuning. A recent survey [55] on TSFMs drew an important distinction between work that pre-trains transformers directly on raw time-series data and work that adapts existing large-language-model (LLM) backbones. Recently, the emergence of several TSFMs illustrate how these architectural ideas are realised at scale:
@@ 2.4 Summary and Research Questions
This study investigates the zero-shot forecasting capabilities of TSFMs in a macroeconomic forecast, deliberately foregoing any domain-specific features. By directly applying out-of-the-box, pre-trained TSFMs, we establish fundamental performance baselines that reveal how effectively these models can leverage their learned representations to forecast economic time series and perform under major shocks, all without any fine-tuning, and utilize national GDP and industry sectors data as representative case study for generalisable research questions.
  • RQ1 How effective are state-of-the-art TSFMs for zero-shot univariate forecasting of macroeconomic and industry-level time series?
  • RQ2 To what extent do TSFMs forecast remain stable when confronted with periods of extreme volatility and significant economic disruption?

### v2

@@ Abstract
This study investigates the transfer learning capabilities of Time Series Foundation Models (TSFMs) under the zero-shot setup, to forecast macroeconomic indicators. New TSFMs are continually emerging, offering significant potential to provide ready-trained and accurate forecasting models that generalise across a wide spectrum of domains. However, the transferability of their learning to many domains, especially economics, is not well understood. To that end, we study TSFM’s performance profile for forecasting economic, bypassing the need for training bespoke econometric models using extensive training datasets. Our experiments were conducted on a univariate case study dataset, in which we rigorously back-tested three state-of-the-art TSFMs (Chronos, TimeGPT and Moirai) under data-scarce conditions and structural breaks. Our results demonstrate that appropriately engineered TSFMs can internalise rich economic dynamics, accommodate regime shifts, and deliver well-behaved uncertainty estimates out of the box, while matching and exceeding state-of-the-art multivariate models currently used on this domain. Our findings suggest that, without any fine-tuning and additional multivariate inputs, TSFMs can match or outperform classical models under both stable and volatile economic conditions. However, like all models, they are vulnerable to performance degradation during periods of rapid shocks, though they recover the forecasting accuracy faster than classical models. The findings offer guidance to practitioners on when zero-shot deployments are viable for macroeconomic monitoring and strategic planning.
Keywords transfer learning, GDP forecasting, time series foundation models, time series forecasting, zero-shot forecasting
@@ 1 Introduction
Macroeconomic indicators, such as gross domestic product (GDP), consumer price inflation, and the unemployment rate, serve as gauges for the economy’s direction. These indicators distil vast amounts of data into manageable signals of aggregate demand, supply-side capacity, and financial conditions. Timely and reliable forecasts of gross domestic product (GDP) are essential inputs for policy, finance, and corporate planning worldwide. Governments embed medium-term growth scenarios in budget frameworks, debt-sustainability analyses, and multi-year spending envelopes [1]. Prudential regulators feed GDP paths into system-wide stress tests to gauge the resilience of banks and insurers under adverse macroeconomic conditions [2]. Investors rebalance portfolios when growth expectations shift, treating GDP as a succinct proxy for business-cycle momentum [3]. Multinational firms align production schedules, inventories, and staffing levels with headline and sectoral GDP projections [4], while credit-rating agencies and development institutions incorporate forward-looking GDP assumptions into sovereign-risk assessments and concessional-finance formulas [5].
@@
However, producing dependable forecasts remains challenging: structural breaks, measurement errors, and sudden shocks can quickly erode model performance, thereby motivating continuous innovation, from classical econometric combinations to modern machine-learning approaches [6]. Forecasting GDP accurately across countries presents significant challenges. Firstly, early data releases are frequently revised, sometimes to the extent of reversing the sign of reported growth. This means models must contend with evolving truths rather than a fixed target [7]. Secondly, structural changes, such as those triggered by commodity-price super-cycles, financial crises, natural disasters, and global pandemics, can abruptly disrupt historical relationships and invalidate previously stable parameters [8, 9, 10, 11]. Lastly, sector-level GDP series display diverse seasonal patterns and react unevenly to external-demand shocks, hindering the transfer of information from one industry to another [12].
@@
For decades, classical time series models, such as autoregressive integrated moving-average (ARIMA) frameworks [13] and vector autoregressions (VARs) [14], have been used as forecasting tools for this domain due to their transparency, tractability, and ease of re-estimation. However, their core assumptions of linearity and stable parameters rarely hold in practice, and sudden shifts in policy regimes and technology shocks can all invalidate coefficients calibrated on historical data, causing forecast accuracy to deteriorate rapidly as the horizon lengthens [15]. Empirical surveys have shown that beyond a few quarters ahead, even well-specified and rich multivariate systems seldom outperform simple persistence or random-walk benchmarks [16]. This challenge is further compounded by the constant moving of the goalposts in the form of data revisions, where preliminary GDP releases and other key indicators are often substantially updated, meaning models trained on early vintages chase a moving target and deliver the least reliability precisely when decision-makers most need clarity [17]. Recent studies have explored more sophisticated modelling approaches such as mixed-frequency factor and Bayesian models, which integrate hundreds of monthly indicators to sharpen nowcasts [18], as well as tree-based ensembles and hybrid neural networks capable of discovering non-linear interactions [19]. However, the more recent modelling approaches are accompanied by high overheads in terms of expertise and computation for modest returns in improved accuracies, while being dependent on provision and access to real-time multivariate inputs [20].
@@
Recent advancements in AI have introduced a new class of forecasting tools, namely Time Series Foundation Models (TSFMs), with the potential to mitigate longstanding challenges in macroeconomic forecasting [21]. TSFMs attempt to leverage transfer learning, which reuses representations learned on a source dataset in order to re-purpose it for making forecasts on a target task. TSFMs are pre-trained on millions of heterogeneous time series with knowledge captured in parameters that in theory can then be transferred to improve performance on a different target either with or without model fine-tuning [22]. Under the zero-shot regime, no further fine-tuning is undertaken, and thus the transferability of TSFMs to accurately forecast future values can then be explored in the purest form, thus significantly reducing the modelling resource overheads [23]. These large pre-trained TSFMs, Nixtla’s TimeGPT, Amazon’s Chronos, and Salesforce’s Moirai treat numeric sequences as language tokens and leverage transformer backbones trained on extensive datasets. TimeGPT offers a “plug-and-play” API capability providing zero-shot forecasts without local fine-tuning, while Moirai extends this paradigm to multivariate forecasting with exogenous covariates. Models like Chronos have been used in literature on domains such as electricity, traffic, and retail data, where they have reported better performances than tuned statistical baselines with minimal feature engineering [24, 25, 26]. Given the largely unexplored capabilities of these models in a macroeconomic context, our study seeks to explore to what degree the current cutting-edge TSFMs can generalise across data vintages, sectors, and structural breaks in economic datasets, under the “out-of-the-box”, zero-shot settings. For our study, we used macroeconomic data from New Zealand as a case study, which presents an unusually demanding testbed, given the vulnerability of its small economy to external factors, as well as its exposure to commodity-price cycles, natural disasters, and external-demand shocks [27], together with the frequent revisions and thus the uncertainty of macroeconomic indicators estimates [28].
@@
While acknowledging that optimal macroeconomic forecasting ideally requires the integration of rich, real-time features capturing key economic indicators as model inputs, this study adopts a distinct approach. We focus specifically on evaluating the pure zero-shot forecasting capabilities of TSFMs demonstrate true zero-shot transfer learning without incorporating such external features or domain-specific covariates. This deliberate choice is made to establish fundamental performance limits and explore the inherent forecasting potential of these models based solely on their pre-training, thereby establishing baselines of their transfer abilities, given the current paucity of research applying TSFMs, particularly in a zero-shot manner, to the economic domain. The contributions and novelty of this work can be summarised as:
  • Empirical benchmark: we provide the zero-shot evaluation of leading TSFMs (Chronos, Moirai, TimeGPT) against classical econometric baselines from the Reserve Bank of New Zealand (RBNZ) forecasts, covering New Zealand’s national GDP and sectoral industries.
  • Performance measurement: we demonstrate that TSFMs outperform other classical methods across various horizons, including RBNZ’s benchmark models, and thus we establish their utility under certain conditions.
  • Operational guidance: we offer actionable insights for policy analysts by mapping the boundary conditions under which zero-shot TSFMs serve as low-maintenance forecasting tools for practitioners or economists. We also identify scenarios where lightweight classical models remain preferable.
2 Related works
@@ 2.1 Forecasting Difficulty for Macroeconomic Indicators
Forecast accuracy for macroeconomic aggregates is fundamentally constrained by low signal-to-noise ratios. A long tradition of forecast evaluation studies shows that, once the horizon stretches beyond the now-cast and the subsequent quarter, point predictions of real GDP growth seldom beat a naive random walk, let alone a purely random-direction guess, when measured against out-of-sample performance [16]. Persistence forecasts, which simply carry forward the latest observed growth rate, provide a standard benchmark over which, even multivariate econometric systems rarely improve upon this baseline after the first step in the forecast horizon [29]. Even median private-sector projections at a four-quarter horizon exhibit RMSEs statistically indistinguishable from persistence [30].
@@
To complement algorithmic advances, researchers now emphasise the timing of data arrival. Mixed-frequency and real-time approaches integrate high-frequency indicators, electronic-card transactions, and daily mobility into quarterly GDP nowcasts [36], providing policymakers near-instant feedback during shocks [37, 38]. Model builders also adapt specifications to the country context. In open economies like New Zealand, global commodity prices, foreign demand, and idiosyncratic domestic cycles jointly shape growth dynamics; assessing how well models internalise these influences remains an active line of inquiry [39].
@@ 2.2 Modern and Emerging Forecasting Approaches
Building on these statistical platforms, institution-specific suites provide operational nowcasts. The Federal Reserve Bank of New York’s medium-scale DSGE integrates theory-consistent shocks with factor information for policy analysis [46]. In contrast, the Atlanta Fed’s GDPNow decomposes each GDP sub-aggregate via bridge equations and updates almost daily, offering a transparent, additive view of U.S. growth [47]. More recently, attention has shifted to machine learning and hybrid methods that relax linearity and exploit high-dimensional features. Ensemble trees (random forests, gradient boosting) already outperform factor and penalised-regression baselines on Dutch GDP nowcasts [45]. Deep neural networks, especially Long Short-Term Memory (LSTM) models, capture nonlinear temporal dependencies and have beaten ARIMA benchmarks in volatile settings [48]. Hybrid ensembles push further by blending economic structure with ML flexibility: weighting forecasts from a time-varying-coefficient DFM and a recurrent neural network by inverse MSE reduces U.S. GDP errors beyond either model alone. At the same time, broader model-averaging strategies remain popular for error reduction [49]. A clear progression from linear autoregressions through factor-based systems and institutional nowcasting dashboards to data-hungry ML hybrids, each stage addressing limitations exposed by the last and setting the stage for evaluating emerging foundation-model approaches.
2.3 Zero-Shot Transfer Learning for Macroeconomic Forecasting
Zero-shot transfer learning applies a pre-trained model to a novel domain or task without requiring additional training or fine-tuning on new data. In the macroeconomic context, a model that has been pre-trained by vast heterogeneous datasets of economic indicators, frequencies, and business regimes to forecast fresh and new series that were never part of the training. Historically, this has been a significant challenge for time series analysis. However, recent research indicates that with appropriate architectures and training methodologies, zero-shot transfer learning is now achievable in this field [50]. This means a model pre-trained on a diverse collection of time series can directly forecast unseen time series without any further adjustments. This approach, also known as zero-shot forecasting, relies on the model’s ability to capture universal temporal patterns that transfer effectively to new series. Suppose the characteristics of the new series are adequately represented by the diverse patterns learned during pre-training. In that case, the model can generate reasonable forecasts without specific training for each new series [51]. Given a new domain that was never seen during training to perform inference without gradient updates [52].
$\hat{\mathbf{y}}_{t+1:t+H}=F_{\theta^{\star}}\bigl{(}\mathbf{y}^{\dagger}_{1:t},\,\mathbf{x}^{\dagger}_{1:t+H}\bigr{)}$ (1)
Equation 1 describes a point of forecast for the next $H$ time steps, using only the past $t$ observations $\mathbf{y}^{\dagger}_{1:t}$ and any covariates $\mathbf{x}^{\dagger}_{1:t+H}$, without any parameter updates, provides the forecast of the distribution over the future $H$ prediction [53].
The feasibility of this approach has significantly increased with the emergence of TSFMs. These are large models trained on heterogeneous time series data from numerous domains. The benefits of zero-shot transfer learning are substantial, it eliminates the need for task-specific training and avoids the requirement for large target datasets. Furthermore, zero-shot forecasting specifically focuses on predicting future values for new time series by leveraging knowledge from a broad pre-trained model, treating each new time series as a zero-shot task. Recent studies have demonstrated surprisingly strong results for zero-shot forecasting using pre-trained models [54].
@@ 2.4 Zero-Shot TSFMs in Economic Forecasting
That early evidence transformers can act as generic sequence learners came from the Frozen Pre-trained Transformer (FPT) experiment, where a language-model backbone was kept entirely frozen across diverse time series tasks, showing that self-attention can operate as a domain computation [55]. Follow-up work systematises this line of research by mapping architectures, pre-training objectives, adaptation strategies, and data modalities [56]. While architectural variants such as encoder–decoder hybrids [57], sparse-attention blocks [58], and decomposition-style residual paths [59] dominate now with networks of millions of parameters that are pre-trained on a vast heterogeneous collection of time series. These TSFMs are typically transformer-based pre-trained models on massive collections of time series data, enabling zero-shot capabilities in forecasting in the concept of zero-shot transfer learning, which allows models to generalise to unseen tasks by leveraging knowledge gained from previously seen data [60].
@@
TSFMs are based on the transformer architecture, shown in Figure 1. TSFMs aim for broad transferability with little or no task-specific fine-tuning. A recent survey [62] on TSFMs drew an important distinction between work that pre-trains transformers directly on raw time series data and work that adapts existing large-language-model (LLM) backbones. Recently, the emergence of several TSFMs illustrates how these architectural ideas are realised at scale:
@@ 2.5 Summary and Research Questions
This study investigates zero-shot forecasting capabilities of TSFMs to achieve zero-shot transfer learning in a macroeconomic forecast, deliberately foregoing any domain-specific features. By directly applying out-of-the-box, pre-trained TSFMs, we establish fundamental performance baselines that reveal how effectively these models can leverage their learned representations to forecast economic time series and perform under major shocks, all without any fine-tuning, and utilise national GDP and industry sectors data as a representative case study for generalisable research questions.
  • RQ1 How effective are state-of-the-art TSFMs for zero-shot univariate forecasting with zero-shot transfer learning of macroeconomic and industry-level time series?
  • RQ2 To what extent do zero-shot TSFMs forecast remain stable when confronted with periods of extreme volatility and significant economic disruption?
@@ 3.1.1 RBNZ Operational Dataset
3.2 Baseline Models
We benchmark TSFMs against four widely used baselines, the persistence model, ARIMA model, the least-squares boosting ensemble (LSBoost), and the factor model.
3.2.1 Persistence model
The persistence forecasting model is given by equation 2.
$\hat{y}*{t+h}=y_{t}\quad\text{for all }h\geq 1,$ (2)
3.2.2 ARIMA model
The Autoregressive–Integrated–Moving-Average (ARIMA) by Nixtla [68] is an automatic process that employs a stepwise search procedure to identify optimal ARIMA and seasonal orders. This is achieved by combining unit-root tests for stationarity with the minimisation of information criteria. Models form a parsimonious yet expressive class for linear dynamics and remain a statistical baseline in forecasting The seasonal form, denoted $\text{ARIMA}(p,d,q)\times(P,D,Q)$, combines non-seasonal and seasonal operators. The model acts as a dynamic statistical benchmark, adapting to each sector’s unique autocorrelation structure and seasonality patterns. It applies differencing to handle trend and drift components and fits candidate models using a maximum search space to maintain computational efficiency.
3.2.3 LSBoost(Least-Squares Boosting)
LSBoost is a gradient boosting method in algorithm 1, which minimises the squared-error (least-squares) loss, effectively performing a functional gradient-descent step in the space of functions in each boosting round. This machine learning ensemble method sequentially combines a finite set of weak learners to improve predictive performance for regression problems [69]. It works by fitting each new learner to the residual errors of the previous ones.
Input: Training set $\{(x_{i},y_{i})\}_{i=1}^{N}$, number of iterations $M$
Output: Boosted predictor $F_{M}(x)$
Define loss function:
$L(y,F)\;=\;\tfrac{(y-F)^{2}}{2}\quad\text{and let }F_{m}(x)\text{ be the current regression model.}$.
[…]
for $m\leftarrow 1$ to $M$ do
   Compute residuals:
$\tilde{y}_{i}\;\leftarrow\;y_{i}\;-\;F_{m-1}(x_{i})\quad(i=1,\dots,N)$
Fit weak learner parameters $(\rho_{m},\alpha_{m})$ by
$(\rho_{m},\alpha_{m})\;=\;\arg\min_{\rho,\alpha}\sum_{i=1}^{N}\bigl{[}\tilde{y}_{i}\;-\;\rho\,h(x_{i};\alpha)\bigr{]}^{2}$
Update model:
$F_{m}(x)\;\leftarrow\;F_{m-1}(x)\;+\;\rho_{m}\,h\bigl{(}x;\alpha_{m}\bigr{)}$
return $F_{M}(x)$
Algorithm 1 LSBoost Algorithm [70]
3.2.4 Factor model
Factor models are a class of statistical models designed to explain the co-movement among a large panel of observed variables $x_{it}$ and a smaller set of latent factors $F_{t}$, represented by equation 3.
$x_{it}=\boldsymbol{\lambda}_{i}^{\top}F_{t}+e_{it},\qquad i=1,\dots,N,\;t=1,\dots,T,$ (3)
where $F_{t}\in\mathbb{R}^{r}$ with $r\ll N$ captures the common dynamics, $\boldsymbol{\lambda}_{i}$ are series-specific, and $e_{it}$ are errors that are allowed weak cross-sectional and serial correlation.
In macroeconomic applications, it is often desirable to model the dynamic evolution of the latent factors. Dynamic factor models (DFMs) augment the static framework by equation 4.
$F_{t}=\Phi_{1}F_{t-1}+\dots+\Phi_{p}F_{t-p}+u_{t},\quad u_{t}\sim\mathcal{N}(0,\Sigma_{u}),$ (4)
which embeds lead-lag relationships among economic indicators. This model strategy underlies modern nowcasting systems that blend hundreds of monthly and high-frequency indicators into real-time GDP estimates [71].
3.3 Time Series Foundation Models
We benchmark each foundation model against baseline models, providing comprehensive comparisons.
3.3.1 TimeGPT-1 model
TimeGPT-1 is a large foundation model for time-series forecasting that adapts the Transformer architecture to temporal data. The model is trained at scale on a massive and diverse dataset of over 100 billion data points, which includes time series from various domains like finance, economics, healthcare, weather, IoT sensors, and web traffic. This extensive and heterogeneous dataset includes series with multiple seasonalities, different cycle lengths, various trends, and significant amounts of noise. Training on such a wide variety of time-series patterns allows TimeGPT-1 to generalise effectively. It can produce accurate zero-shot forecasts on entirely new time series without any specific re-training. For instance, the model can accurately forecast New Zealand’s GDP and its industry sub-series without any fine-tuning or parameter adjustments. Architecturally, the model uses an encoder-decoder Transformer with stacked self-attention blocks, residual connections, layer normalisation and positional encodings for the input window. A final projection layer maps the decoder’s output to the desired forecast horizon, and the learned parameters are frozen to maintain its zero-shot forecasting capability [24]. This is demonstrated by direct comparison to baseline models.
3.3.2 Chronos model
Chronos is a pre-trained time series forecasting model that re-frames forecasting as a language modelling task. The model is trained on a large corpus of publicly available time series data from diverse sectors like energy, finance, and retail. To enable this, Chronos first converts real-valued time series observations into a sequence of discrete tokens by using scaling and quantisation. It trains a transformer model (specifically, an off-the-shelf T5-style encoder-decoder) on these tokens using cross-entropy loss. This approach allows Chronos to learn generalised temporal representations that transfer across different datasets, resulting in enhanced forecasting accuracy. Forecasting proceeds by taking the final hidden states of the model for a given token sequence and, using a conditional distribution for the next token, autoregressively sampling new tokens. These tokens are then de-quantised to produce the zero-shot forecasting [26] to compare to baseline models.
3.3.3 Moirai model
Moirai, a masked-encoder universal Transformer, was developed by Salesforce AI Research for zero-shot forecasting. It operates across diverse time-series domains using the Large-scale Open Time-Series Archive (LOTSA), which includes datasets from nine domains: energy, transport, climate, cloud operations, web, sales, nature, economics/finance, and healthcare. To handle high-frequency series efficiently and reduce the quadratic cost of self-attention, the model uses larger patch sizes for embedding. It also introduces multiple patch-size input and output projection layers. The model employs an any-variate attention mechanism, which flattens multivariate time series into a single sequence. This mechanism utilises rotary position embeddings for time indices and learned binary attention biases for variate indices, optimising the negative log-likelihood by using the parameters of a mixture of distributions. Moirai is available in three sizes: Moirai-Small (6 layers, 14M parameters), Moirai-Base (12 layers, 91M parameters), and Moirai-Large (24 layers, 311M parameters). Moirai provides competitive forecasts compared with fully pre-trained models, demonstrating the ability to transfer across domains and frequencies, using a single set of trained parameters to handle any domain and frequency in a zero-shot manner without fine-tuning or re-training. This approach allows Moirai to deliver flexible and accurate zero-shot forecasting on out-of-distribution datasets, handling diverse data from various sectors and exogenous covariates [25] for direct evaluation against baseline models.